OpenStudy (dengeki_daisy):

What is the area of a sector with a central angle of 160 degrees and a diameter of 5.8 m? Round the answer to the nearest tenth. a) 8.1 m² b) 4.0 m² c) 47.0 m² d) 11.7 m²

OpenStudy (welshfella):

area = (160/360) pi r^2 where r = 1/2 diameter = 1/2 * 5.8 = 2.9

OpenStudy (dengeki_daisy):

so 4.0

OpenStudy (dengeki_daisy):

welshfella

OpenStudy (dengeki_daisy):

can someone help

OpenStudy (debbieg):

Did you use the sector area formula that welshfella gave you above? Sector area is basically just a proportion of a total circle area (which is pi r^2)

OpenStudy (debbieg):

Area of a circle = pi * r^2 Area of a sector = area of the total circle x the proportion of the total circle that the sector is If angle is given in degrees, then that proportion is (angle degrees / 360)

OpenStudy (welshfella):

no not 4.0 plug the expression into ur calculator (160/360) * 3.14 * 2.9^2 - use pi = 3.14 will be OK

OpenStudy (dengeki_daisy):

I did and I got 11.7. Is that right?

OpenStudy (dengeki_daisy):

welshfella

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

gotta go